# Learning Humanoid Standing-up Control across Diverse Postures

### v1

@@ Abstract
Standing-up control is crucial for humanoid robots, with the potential for integration into current locomotion and loco-manipulation systems, such as fall recovery. Existing approaches are either limited to simulations that overlook hardware constraints or rely on predefined ground-specific motion trajectories, failing to enable standing up across postures in real-world scenes. To bridge this gap, we present HoST (Humanoid Standing-up Control), a reinforcement learning framework that learns standing-up control from scratch, enabling robust sim-to-real transfer across diverse postures. HoST effectively learns posture-adaptive motions by leveraging a multi-critic architecture and curriculum-based training on diverse simulated terrains. To ensure successful real-world deployment, we constrain the motion with smoothness regularization and implicit motion speed bound to alleviate oscillatory and violent motions on physical hardware, respectively. After simulation-based training, the learned control policies are directly deployed on the Unitree G1 humanoid robot. Our experimental results demonstrate that the controllers achieve smooth, stable, and robust standing-up motions across a wide range of laboratory and outdoor environments (Fig. 1). Videos are available on our project page.
@@ I Introduction
Reinforcement learning (RL) offers an alternative effective framework for humanoid locomotion and whole-body control [36, 13, 4, 53], benefiting from minimal modeling assumptions. However, compared to these tasks that partially decouple upper- and lower-body dynamics, RL-based standing-up control involves a highly dynamic and synergistic maneuver on both halves of the body. This complex maneuver features time-varying contact points [17], multi-stage motor skills [29], and precise angular momentum control [11], making RL exploration challenging. Although predefined motion trajectories can guide RL exploration, they are typically limited to ground-specific postures [35, 36, 51, 12], leaving the scalability to other postures unclear. Conversely, training RL agents from scratch with wide explorative strategies on the ground can lead to violent and abrupt motions that hinder real-world deployment [46], particularly for robots with many actuators and wide joint limits. In summary, learning posture-adaptive, real-world deployable standing-up control with RL remains an open problem (see Table I).
@@ III Problem Formulation
III-1 State Space
We hypothesize that the proprioceptive states of robots provide sufficient information for standing-up control in our target environments. We thus include the proprioceptive information read from robot’s Inertial Measurement Unit (IMU) and joint encoders into the state $s_{t}=[\omega_{t},r_{t},p_{t},\dot{p}_{t},a_{t-1},\beta]$, where $\omega_{t}$ is the angular velocity of robot base, $r_{t}$ and $p_{t}$ are the roll and pitch, $p_{t}$ and $\dot{p}_{t}$ are positions and velocities of the joints, $a_{t-1}$ is the last action, and $\beta\in(0,1]$ is a scalar that scale the output action. Given the contact-rich nature of the standing-up task, we implicitly enhance contact detection by feeding the policy with the previous five states [15].
@@ IV-A2 Exploration Challenges (Section IV-C)
Despite multi-critic RL, exploration remains difficult due to the robot’s high degrees of freedom and wide joint limits. Drawing inspiration from human infant skill development [6], we facilitate exploration by applying a curriculum-based vertical pulling force.
@@ IV-A4 Sim-to-Real Gap (Section IV-E)
A significant challenge is the sim-to-real gap. We address this through two strategies: (1) designing diverse terrains to better simulate real-world starting postures, and (2) applying domain randomization [48] to reduce the influence of physical discrepancies between simulation and real world.
@@ IV-B Reward Functions & Multiple Critics
Considering the multi-stage nature of the task, we divide the task into three stages: righting the body $h_{\mathrm{base}}<H_{\mathrm{stage1}}$, rising the body $h_{\mathrm{base}}>H_{\mathrm{stage2}}$, and standing $h_{\mathrm{base}}>H_{\mathrm{stage2}}$, indicated by the height of the robot base $h_{\mathrm{base}}$. Corresponding reward functions are activated at each stage. We then classify reward functions into four groups: (1) task reward $r^{\mathrm{task}}$ that specifies the high-level task objectives, (2) style reward $r^{\mathrm{style}}$ that shapes the style of standing-up motion, (3) regularization reward $r^{\mathrm{regu}}$ that further regularizes the motionw, and (4) post-task reward $r^{\mathrm{post}}$ that specify the desired behaviors after successful standing up, i.e., stay standing. The overall reward function is expressed as follows:
@@ IV-C Force Curriculum as Exploration Strategy
The primary exploration challenges emerge during the transition from falling to stable kneeling, a stage that proves difficult to explore effectively through random action noise alone. While human infants are likely to learn motor skills with external supports [6], it inspires us to design environmental assistance to accelerate the exploration. Specifically, we apply an upward force $\mathcal{F}$ on the robot base, which is largely set at the start of training. This force takes effect only when the robot’s trunk achieves a near-vertical orientation, indicating a successful ground-sitting posture. The force magnitude decreases progressively as the robot can maintain a target height at the end of the episode. See more details in Section -A.
IV-D Motion Smoothness
@@ IV-D1 Action Bound (Rescaler)
Humanoid robots often feature many DoFs, each equipped with wide position limits and high-power actuators. This configuration often results in violent motions after RL training, characterized by violent ground hitting and rapid bouncing movements. While setting low action bounds could mitigate this behavior, it might prevent the robot from exploring effective standing-up motions. To this end, we introduce an action rescaler $\beta$ to scale the action output, implicitly controlling the bound of the maximal torques on each actuator. This scale coefficient gradually decreases like vertical force reduction. See more details in Section -A.
@@
Base CoM offset $\mathcal{U}(-0.03,0.03)$m
@@ VIII Limitations and Future Directions
Integration with existing humanoid systems. Although integration with existing humanoid systems is not demonstrated in this paper, we envision that standing-up control can be effectively incorporated into current humanoid frameworks to extend real-world applications.

### v2

@@ Abstract
Standing-up control is crucial for humanoid robots, with the potential for integration into current locomotion and loco-manipulation systems, such as fall recovery. Existing approaches are either limited to simulations that overlook hardware constraints or rely on predefined ground-specific motion trajectories, failing to enable standing up across postures in real-world scenes. To bridge this gap, we present HoST (Humanoid Standing-up Control), a reinforcement learning framework that learns standing-up control from scratch, enabling robust sim-to-real transfer across diverse postures. HoST effectively learns posture-adaptive motions by leveraging a multi-critic architecture and curriculum-based training on diverse simulated terrains. To ensure successful real-world deployment, we constrain the motion with smoothness regularization and implicit motion speed bound to alleviate oscillatory and violent motions on physical hardware, respectively. After simulation-based training, the learned control policies are directly deployed on the Unitree G1 humanoid robot. Our experimental results demonstrate that the controllers achieve smooth, stable, and robust standing-up motions across a wide range of laboratory and outdoor environments (Fig. 1). Videos and code are available on our project page.
@@ I Introduction
Reinforcement learning (RL) offers an alternative effective framework for humanoid locomotion and whole-body control [36, 13, 4, 54], benefiting from minimal modeling assumptions. However, compared to these tasks that partially decouple upper- and lower-body dynamics, RL-based standing-up control involves a highly dynamic and synergistic maneuver on both halves of the body. This complex maneuver features time-varying contact points [17], multi-stage motor skills [29], and precise angular momentum control [11], making RL exploration challenging. Although predefined motion trajectories can guide RL exploration, they are typically limited to ground-specific postures [35, 36, 52, 12], leaving the scalability to other postures unclear. Conversely, training RL agents from scratch with wide explorative strategies on the ground can lead to violent and abrupt motions that hinder real-world deployment [46], particularly for robots with many actuators and wide joint limits. In summary, learning posture-adaptive, real-world deployable standing-up control with RL remains an open problem (see Section -B).
@@ III Problem Formulation
III-1 Observation Space
We hypothesize that the proprioceptive states of robots provide sufficient information for standing-up control in our target environments. We thus include the proprioceptive information read from robot’s Inertial Measurement Unit (IMU) and joint encoders into the state $s_{t}=[\omega_{t},r_{t},q_{t},p_{t},\dot{p}_{t},a_{t-1},\beta]$, where $\omega_{t}$ is the angular velocity of robot base, $r_{t}$ and $q_{t}$ are the roll and pitch, $p_{t}$ and $\dot{p}_{t}$ are positions and velocities of the joints, $a_{t-1}$ is the last action, and $\beta\in(0,1]$ is a scalar that scale the output action. Given the contact-rich nature of the standing-up task, we implicitly enhance contact detection by feeding the policy with the previous five states [15].
@@ IV-A2 Exploration Challenges (Section IV-C)
Despite multi-critic RL, exploration remains difficult due to the robot’s high degrees of freedom and wide joint limits. Drawing inspiration from human infant skill development [6, 49], we facilitate exploration by applying a curriculum-based vertical pulling force on the base link of the robot.
@@ IV-A4 Sim-to-Real Gap (Section IV-E)
A significant challenge is the sim-to-real gap. We address this through two strategies: (1) designing diverse terrains to better simulate real-world starting postures, and (2) applying domain randomization [48] to reduce the influence of physical discrepancies between simulation and the real world.
@@ IV-B Reward Functions & Multiple Critics
Considering the multi-stage nature of the task, we divide the task into three stages: righting the body $h_{\mathrm{base}}<H_{\mathrm{stage1}}$, rising the body $H_{\mathrm{stage1}}<h_{\mathrm{base}}<H_{\mathrm{stage2}}$, and standing $h_{\mathrm{base}}>H_{\mathrm{stage2}}$, indicated by the height of the robot base $h_{\mathrm{base}}$. Corresponding reward functions are activated at each stage. We then classify reward functions into four groups: (1) task reward $r^{\mathrm{task}}$ that specifies the high-level task objectives, (2) style reward $r^{\mathrm{style}}$ that shapes the style of standing-up motion, (3) regularization reward $r^{\mathrm{regu}}$ that further regularizes the motionw, and (4) post-task reward $r^{\mathrm{post}}$ that specify the desired behaviors after successful standing up, i.e., stay standing. The overall reward function is expressed as follows:
@@ IV-C Force Curriculum as Exploration Strategy
The primary exploration challenges emerge during the transition from falling to stable kneeling, a stage that proves difficult to explore effectively through random action noise alone. While human infants are likely to learn motor skills with external supports [6, 49], it inspires us to design environmental assistance to accelerate the exploration. Specifically, we apply an upward force $\mathcal{F}$ on the robot base, which is largely set at the start of training. This force takes effect only when the robot’s trunk achieves a near-vertical orientation, indicating a successful ground-sitting posture. The force magnitude decreases progressively as the robot can maintain a target height at the end of the episode. See more details in Section -B.
IV-D Motion Constraints
@@ IV-D1 Action Bound (Rescaler)
Humanoid robots often feature many DoFs, each equipped with wide position limits and high-power actuators. This configuration often results in violent motions after RL training, characterized by violent ground hitting and rapid bouncing movements. While setting low action bounds could mitigate this behavior, it might prevent the robot from exploring effective standing-up motions. To this end, we introduce an action rescaler $\beta$ to scale the action output, implicitly controlling the bound of the maximal torques on each actuator. This scale coefficient gradually decreases like vertical force reduction. See more details in Section -B.
@@
Base CoM offset $\mathcal{U}(-d,d)$m, $d=0.12(XY),0.08(Z)$
@@ VI-C Emergent Properties
Figure 12: More diverse postures. HoST can learn across prone postures on the ground. The learn policies can also handle side-lying postures.
@@
VI-D Prone and Side-lying Postures on the Ground
We demonstrate that HoST is capable of learning across prone postures on the ground, as visualized in Fig. 12. Besides, the learned policies can also handle side-lying postures without any tuning. However, there are significant differences in motion patterns between prone and supine postures. This somehow limits our method: when training from posture postures, harder constraints on hip joints are necessary to prevent violent motions, making the feasibility of joint training from prone and supine postures unclear currently.
VI-E Extend HoST to Large-size Humanoid Robots
Figure 13: Extension to large-size robots. HoST can be easily extended to Unitree H1 and H1-2 humanoid robots with minor hyperparameter tuning.
We believe that standing-up control is more challenging in larger humanoid robots than in G1 due to their increased weight and the limited actuators. As an initial test, we extend HoST to Unitree H1 and H1-2. Simulation and real-world motions are visualized in Fig. 13. Compared to G1, we observe greater reliance on (i) upper-body contact with the ground and (ii) high hip actuation. While successful, two sim-to-real gaps emerge: (i) the need for high-stiffness joints to compensate for insufficient torques and, which is consistent with the observation in Section VI-B; (ii) noticeable deviations in upper-body posture. It remains unclear whether these gaps originate from our framework or hardware limitations. Identifying the source of these gaps is valuable in the future.
@@ VIII Limitations and Future Directions
Integration with existing humanoid systems. Although this paper does not demonstrate integration with existing humanoid systems, we envision that standing-up control can be effectively incorporated into current humanoid frameworks to extend their real-world applications.
Acknowledgments
This work is funded in part by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory.